# Modified Abelian and SU(2) Wilson theories on a lattice from a non-compact regularization

### v1

@@ III.2 Modified half-compact Lagrangian
Again one can look at the difference between the present non-compact regularization and the half-compact one performing an expansion of the interaction term to leading order in the lattice spacing. The result is found inserting Eq.(54) in Eq.(44)
@@
where we took into account that $\beta=1/g^{2}$. The Lagrangians of the matter fields get obviously the same corrections as in the non-compact regularization.
IV Non-compact SU(2) gauge theories on a lattice
The formulation of non-compact SU(2) gauge theories has been treated in detail [Becc1, ; Becc2, ], and most of the results of this Section about the modification of the compact one can simply be derived from the Abelian case introducing color indices and the relative trace in the proper places. Nevertheless we will report more equations than strictly necessary to make this Section somewhat selfcontained for a reader interested in the SU(2) case only.
@@ IV Non-compact SU(2) gauge theories on a lattice
For a field $\phi$ transforming as in Eq. (9) one can define right/left covariant derivatives and the symmetric covariant derivative as in Eqs. (II.1), (11), resp. They all transform according to Eq. (9).
@@
where $\beta=4/g^{2}$ and Tr is the trace over colours.
Also with SU(2), in order to avoid contributions of $W_{\mu}\approx 1/a$ for which the covariant derivative does not contain an ordinary derivative, one can include in the Lagrangian a potential which must necessarily be a function of the invariant
@@
Notice that $D_{\mu}^{\dagger}D_{\mu}$ is not a quaternion but it is a real function of the lattice sites as in U(1), so that the potential has exactly the form (14), which gives to the auxiliary field $W_{\mu}$ the mass $4\gamma^{2}/a^{2}$. Moreover it damps all the gauge fields, including their zero-modes, so making unnecessary a gauge fixing (see Appendix A for details).
@@ IV.2 Modified compact SU(2) Lagrangian
At variance with the Abelian case, the couplings $\beta,\gamma$ are not independent from each other. Due to asymptotic freedom perturbation theory is justified, and to one loop it gives [Becc2, ]
@@
where $\gamma_{1},\gamma_{2}$ are independent of $\beta$, $\gamma_{1}$ is arbitrary and $\gamma_{2}$ can be, but it has not been evaluated. This result has been confirmed by a calculation in the Hamiltonian formalism [Bora, ] (see Appendix B for details and comparison with numerical simulations). In conclusion, we cannot vary $\gamma$ arbitrarily, but in the scaling region it is certainly large so that the field $s_{\mu}$ is coupled linearly to the Wilson variables as in the Abelian case, and it can be integrated out giving the modified gauge fields Lagrangian
$\mathcal{L}_{\rm G}=\mathcal{L}_{\rm G}^{\rm W}-\frac{\beta^{2}\,a^{4}}{8\gamma^{2}}\,\sum_{\mu}\left[{\rm Tr}\sum_{\nu}U_{P\mu\nu}(x)\right]^{2}\,.$ (77)
Also the evaluation of matter fields Lagrangians proceeds in the same way as in the Abelian case. The covariant derivative of a scalar with flavour quantum number $f$ and colour quantum number $c$ in the fundamental representation for small lattice spacing reads
@@
V Three-dimensional case
@@
Scalar and fermion fields have dimension 1/2, 1 respectively so that their Lagrangians remain unchanged as well.
@@
VI Conclusions
In recent works U(1) models in three dimensions were compared in compact [Bona1, ] and half-compact form [higher-charge, ] getting quite different phase diagrams, while SU(2) models (for which a half-compact formulation does not exist) were studied only in compact form. It seems therefore interesting to extend the comparison to an entirely non-compact lattice formulation which was defined for general non-Abelian theories. Such regularization showed for SU(2) a faster convergence to the continuum, but the price to be paid is the introduction of one auxiliary gauge invariant field. For a first investigation one can estimate the corrections to the Wilson formulation for large mass of the auxiliary field, can be integrated out yielding a Lagrangian expressed in terms of compact gauge fields and matter fields only. For a first estimate one might just evaluate the contribution of the correction at the phase transition.
@@ VI Conclusions
The present result can be extended to QCD, but in this case there are colour auxiliary fields and $D^{\dagger}_{\mu}D_{\mu}$ is no longer a gauge invariant function of the sites but it has an expression involving the generators of the group. We hope to be able to tackle this problem in a future work.
Appendix A Boundary conditions
@@
Boundary conditions however cannot be chosen on the basis of mere calculational convenience, because they may have physical effects; quite generally different boundary conditions favour or disfavour different phases [Camp, ] or select different physical systems as illustrated by the following examples.
Lüscher constructed an effective action for continuous QCD in 4 dimensions with periodic boundary conditions in terms of zero-momentum modes (gauge fields depending on time but constant in space) integrating out all the other modes [Lusc, ]. The mass spectrum so obtained differs from that derived [Gonz, ] assuming the twisted boundary conditions (which forbid zero-momentum modes) firstly studied by ’t Hooft ['tHo, ]. Indeed the latter boundary conditions force fluxes of magnetic and electric fields in the quantization box, defining a physical system different from the one studied by Lüscher.
@@
Appendix B Perturbative calculations versus numerical simulations
@@
  1. a)
    the choice of the partition function (73) or (74);
  2. b)
@@
Concerning point a), in both perturbative calculations [Becc2, ; Bora, ] the partition function (74) was adopted with the approximation $\gamma\gg 1$ in the lower bound of integration over $s_{\mu}$. Therefore to be compared with the quoted perturbative results the numerical simulations should have $\gamma\gg 1$ in the scaling window.
@@
Concerning point b), the arbitrary polynomial has been chosen $\mathcal{P}=-\mathcal{L}_{\rho}$ in the perturbative calculation [Bora, ] and in the simulation [Poli1, ], while it has been chosen $\mathcal{P}=-\mathcal{L}_{\rho}-\mathcal{L}_{l}$ in the perturbative calculation [Becc2, ] and in the simulation [DiCa, ].

### v2

@@ III.2 Modified half-compact Lagrangian
Again one can look at the difference between the present noncompact regularization and the half-compact one performing an expansion of the interaction term to leading order in the lattice spacing. The result is found inserting Eq.(54) in Eq.(44)
@@
where we took into account that $\beta=1/g^{2}$. The Lagrangians of the matter fields get obviously the same corrections as in the noncompact regularization.
IV Noncompact SU(2) gauge theories on a lattice
The formulation of noncompact SU(2) gauge theories has been treated in detail [Becc1, ; Becc2, ], and most of the results of this section about the modification of the compact one can simply be derived from the Abelian case introducing color indices and the relative trace in the proper places. Nevertheless we will report more equations than strictly necessary to make this section somewhat self-contained for a reader interested in the SU(2) case only.
@@ IV Noncompact SU(2) gauge theories on a lattice
For a field $\phi$ transforming as in Eq. (9) one can define right/left covariant derivatives and the symmetric covariant derivative as in Eqs. (II.1), (11), respectively. They all transform according to Eq. (9).
@@
where $\beta=4/g^{2}$ and Tr is the trace over colors.
Also with SU(2), in order to avoid contributions of $W_{\mu}\approx 1/a$ for which the covariant derivative does not contain an ordinary derivative, one can include in the Lagrangian a potential that must necessarily be a function of the invariant
@@
Notice that $D_{\mu}^{\dagger}D_{\mu}$ is not a quaternion but it is a real function of the lattice sites as in U(1), so that the potential has exactly the form (14), which gives to the auxiliary field $W_{\mu}$ the mass $4\gamma^{2}/a^{2}$. Moreover it damps all the gauge fields, including their zero modes, so making unnecessary a gauge fixing (see Appendix IX for details).
@@ IV.2 Modified compact SU(2) Lagrangian
At variance with the Abelian case, the couplings $\beta,\gamma$ are not independent from each other. Due to asymptotic freedom, perturbation theory is justified and to one loop it gives [Becc2, ]
@@
where $\gamma_{1},\gamma_{2}$ are independent of $\beta$, $\gamma_{1}$ is arbitrary and $\gamma_{2}$ can be, but it has not been evaluated. This result has been confirmed by a calculation in the Hamiltonian formalism [Bora, ] (see Appendix X for details and comparison with numerical simulations). In conclusion, we cannot vary $\gamma$ arbitrarily, but in the scaling region it is certainly large so that the field $s_{\mu}$ is coupled linearly to the Wilson variables as in the Abelian case, and it can be integrated out giving the modified gauge fields Lagrangian
$\mathcal{L}_{\rm G}=\mathcal{L}_{\rm G}^{\rm W}-\frac{\beta^{2}\,a^{4}}{8\gamma^{2}}\,\sum_{\mu}\left[{\rm Tr}\sum_{\nu}U_{P\mu\nu}(x)\right]^{2}$ (77)
which contains a negative definite contribution. In the deep continuum limit we can use Eq. (76), getting
$\mathcal{L}_{\rm G}=\mathcal{L}_{\rm G}^{\rm W}-\frac{a^{4}}{8\gamma_{1}^{2}}\,\sum_{\mu}\left[{\rm Tr}\sum_{\nu}U_{P\mu\nu}(x)\right]^{2}\,.$ (78)
In the scaling windows of Ref. [DiCa, ], instead,
[…]
Also the evaluation of matter fields Lagrangians proceeds in the same way as in the Abelian case. The covariant derivative of a scalar with flavor quantum number $f$ and color quantum number $c$ in the fundamental representation for small lattice spacing reads
@@
V Comparison with some previous modifications of Wilson action
Previous improvements of the Wilson action essentially aim at reducing the lattice artifacts. A thorough review of the huge literature on the subject is outside the scope of the present work, and we will restrict ourselves to a short discussion of the reduction of lattice effects close to the continuum and effects originated by large fluctuations of the plaquette.
To the first category belong Wilson’s renormalization group [Wils2, ] and Symanzik’s approach based on a perturbative expansion with respect to the lattice spacing [Syma, ]. Wilson’s approach is based on the fact that under a renormalization group transformation the lattice action moves toward a fixed point along a definite trajectory, without change of the partition function, that is the same as that of the original system and describes the same long range physics reducing the lattice artifacts [LusWei, ]. Combining analytical with numerical calculations one can get “perfect” lattice actions completely cutoff independent even on coarse grained lattices [HasNie, ].
Symanzik’s approach is based on the idea of adding extra terms to Wilson’s action chosen in a way to cancel the lattice artifacts of $O(a^{n})$ to the maximum practically possible order [Syma, ,Wei, ]. Both methods along with some subsequent developments were reviewed by Niedermayer [Nied, ].
Concerning large fluctuations of the plaquette, in order to reduce them two different kinds of actions have been proposed in Refs. [Bane, ] and [Lusc1, ], respectively
$S_{p}=\beta w+\gamma w^{q}\,,$ (83)
with $w$ as a plaquette action,
${\cal L}_{\mu\nu}=\left\{\begin{array}[]{l}\displaystyle{\frac{F^{2}_{\mu\nu}}{1-\frac{1}{\epsilon^{2}}\,F^{2}_{\mu\nu}}}\qquad{\mathrm{if}}\qquad|F_{\mu\nu}|\,<\,\epsilon\\ \\ \infty\qquad{\mathrm{otherwise}}.\end{array}\right.$
Actually the original purpose of the second one was to formulate a gauge-invariant Abelian chiral theory on the lattice and the cutoff of large plaquette fluctuations was instrumental to perform some necessary formal manipulations, but it was later used as a way to reduce plaquette fluctuations [BieFu, ].
We notice that for small lattice spacing and plaquette value the last action becomes
${\cal L}_{\mu\nu}\approx(F_{\mu\nu})^{2}\,\left[1+\frac{1}{\epsilon^{2}}\,F_{\mu\nu}\right]^{2}$ (84)
which is (after the necessary correspondence of parameters is established) identical to ours, Eq. (44), but with opposite sign of the correction. Also in the first one the correction has opposite sign with respect to our expression (44).
Our results have been obtained in a spirit completely different from all the above: the heuristic idea is that starting with noncompact fields one is closer to the continuum since the beginning as confirmed by Ref. [DiCa, ]. From the technical point of view we remind that our correction has been obtained by an expansion in inverse powers of the mass of the auxiliary field and it turns out to be $O(a^{4})$. Our results suggest that the actions in Eqs. (83),(V) should be modified by adding terms correcting for lattice artifacts at small lattice spacing.
We can observe that the present improvement is very simple and does not need evaluation of other terms than the plaquette. Moreover, we think that our improved action might be more convenient than Wilson’s one as a starting point in all the approaches discussed above.
VI Three-dimensional case
@@
Scalar and fermion fields have dimension 1/2, 1, respectively so that their Lagrangians remain unchanged as well.
@@
VII Conclusions
In recent works U(1) models in three dimensions were compared in compact [Bona1, ] and half-compact form [higher-charge, ] getting quite different phase diagrams, while SU(2) models (for which a half-compact formulation does not exist) were studied only in compact form. It seems therefore interesting to extend the comparison to an entirely noncompact lattice formulation that was defined for general non-Abelian theories. Such regularization showed for SU(2) a faster convergence to the continuum, but the price to be paid is the introduction of one auxiliary gauge invariant field.
For a first investigation one can estimate the corrections to the Wilson Lagrangian for large mass of the auxiliary field, which can therefore be integrated out yielding a Lagrangian expressed in terms of compact gauge fields and matter fields only. For a further simplification one can evaluate the contribution of the correction only at the phase transition.
@@ VII Conclusions
The present result can be extended to QCD, but in such a case there are color auxiliary fields and $D^{\dagger}_{\mu}D_{\mu}$ is no longer a gauge invariant function of the sites but it has an expression involving the generators of the group. We hope to be able to tackle this problem in a future work.
VIII Acknowledgments
IX Appendix A: Boundary conditions
@@
Boundary conditions however cannot be chosen on the basis of mere calculational convenience, because they may have physical effects; quite generally different boundary conditions favor or disfavor different phases [Camp, ] or select different physical systems as illustrated by the following examples.
Lüscher constructed an effective action for continuous QCD in four dimensions with periodic boundary conditions in terms of zero-momentum modes (gauge fields depending on time but constant in space) integrating out all the other modes [Lusc, ]. The mass spectrum so obtained differs from that derived [Gonz, ] assuming the twisted boundary conditions (which forbid zero-momentum modes) firstly studied by ’t Hooft ['tHo, ]. Indeed the latter boundary conditions force fluxes of magnetic and electric fields in the quantization box, defining a physical system different from the one studied by Lüscher.
@@
X Appendix B: Perturbative calculations versus numerical simulations
@@
  1. 1.
    The choice of the partition function (73) or (74).
  2. 2.
@@
Concerning point 1, in both perturbative calculations [Becc2, ; Bora, ] the partition function (74) was adopted with the approximation $\gamma\gg 1$ in the lower bound of integration over $s_{\mu}$. Therefore to be compared with the quoted perturbative results the numerical simulations should have $\gamma\gg 1$ in the scaling window.
@@
Concerning point 2, the arbitrary polynomial has been chosen $\mathcal{P}=-\mathcal{L}_{\rho}$ in the perturbative calculation [Bora, ] and in the simulation [Poli1, ], while it has been chosen $\mathcal{P}=-\mathcal{L}_{\rho}-\mathcal{L}_{l}$ in the perturbative calculation [Becc2, ] and in the simulation [DiCa, ].